How many liters of 20% sliver must be added to 80 liters of 50% sliver to make an alloy that is 40% sliver ?
Found 2 solutions by josgarithmetic, macston:
Answer by josgarithmetic(39618)

Best thing is to understand how these mixture problems work.
Check and study this lesson: Lesson about two part mixtures to find one quantity when the other quantity is known.

Answer by macston(5194)

.
T=amount of 20%; F=amount of 50%
.
0.20T+0.50F=0.40(T+F)
0.20T+0.50(80)=0.40(T+80)
0.20T+40.0=0.40T+32.0
8.0=0.20T
40=T
.
ANSWER: 40 liters of 20% solution should be added.
.